# Multi-Pass Streaming Lower Bounds for Uniformity Testing

### v1

@@ 3 Hybrid Methods
We poissonize the sample size and consider streams of random length $\mathrm{Poi}(n)$. This is essentially equivalent to working with exactly $n$ samples. Let $D_{\mathrm{unif}}$ denote a stream of $\mathrm{Poi}(n)$ uniform samples from $[2m]$, and $D_{\mathrm{bias}}$ a stream of $\mathrm{Poi}(n)$ elements drawn from the planted bias distribution. We aim to upper bound the achievable advantage on distinguishing between $D_{\mathrm{unif}}$ and $D_{\mathrm{bias}}$.
@@ 3.1 Hybrid Distributions
Suppose $k\in[n]$. Let $D^{k,\frac{n}{k}}$ be a distribution over sequences of elements. To sample from $D^{k,\frac{n}{k}}$. We first draw $k$ sequences $X^{1},\dots,X^{k}$ as follows. For each $i\in[k]$:
@@
    then draw $X^{i}$ as a sequence of $\mathrm{Poi}(n/k)$ independent samples from $P_{Y^{i}}$.
@@
Now, observe that $D_{\mathrm{bias}}\equiv D^{1,n}$. We also have $d_{TV}(D_{\mathrm{unif}},D^{n,1})\leq O(\frac{1}{n})$. To see this, note that $D^{n,1}$ can be understood as the concatenation of $n$ sub-sequences, each with fresh bias pattern $P_{Y^{i}}$. $D^{n,1}$ would “agree” with $D_{\mathrm{unif}}$ if none of the subsequences has length greater than $1$, which happens with probability $O(\frac{1}{n})$ by properties of Poisson distributions and a simple union bound.
Therefore, up to an additive factor of $O(\frac{1}{n})$, the achievable advantage between $D_{\mathrm{unif}}$ and $D_{\mathrm{bias}}$ and that between $D^{n,1}$ and $D^{1,n}$ are the same. In the following, we switch to analyzing $D^{n,1}$ versus $D^{1,n}$.
@@ 3.2 The Hybrid Method
$\mathrm{Adv}(D_{\mathrm{unif}},D_{\mathrm{bias}})\pm O(\frac{1}{n})$ $=\mathrm{Adv}\left(D^{n,1},D^{1,n}\right)$
@@
In order to prove Theorem˜3.3, it suffices to consider its corresponding two-player communication problem, called Hidden Sign Problem.
Definition 3.4 (Hidden Sign Problem).
@@
The rest of the paper will focus on proving a communication lower bound for the Hidden Sign Problem. That is, we will prove the following theorem.
@@
We note that Theorem˜3.5 implies Theorem˜3.3, and hence the main result of our paper.
4 Proof of Theorem˜3.5
In this section, we prove Theorem˜3.5. This is achieved via a series of reductions.
@@ 4.1 The Hidden Index Problem
Finally, in order to analyze the game from Definition˜3.4, we will have to work with a variant of Noisy HIP where we observe not bits but some other forms of signal emitted from two similar sources.
@@ 4.3 Concluding the Proof
We are now ready to analyze the game from Definition˜3.4. We state the following theorem.

### v2

@@ 3 Hybrid Methods
Let $D_{\mathrm{unif}}$ denote a stream of $n$ uniform samples from $[2m]$, and $D_{\mathrm{bias}}$ a stream of $n$ elements drawn from the planted bias distribution. We aim to upper bound the achievable advantage on distinguishing between $D_{\mathrm{unif}}$ and $D_{\mathrm{bias}}$.
@@ 3.1 Hybrid Distributions
Suppose $k\in[n]$ and $k$ divides $n$. Let $D^{k,\frac{n}{k}}$ be a distribution over sequences of elements. To sample from $D^{k,\frac{n}{k}}$, we first draw $k$ sequences $X^{1},\dots,X^{k}$ as follows. For each $i\in[k]$:
@@
    then draw $X^{i}$ as a sequence of $\frac{n}{k}$ independent samples from $P_{Y^{i}}$.
@@
It is easy to observe that $D_{\mathrm{bias}}\equiv D^{1,n}$ and $D_{\mathrm{unif}}\equiv D^{n,1}$. In the following, we analyze the distinguishing advantage between $D^{1,n}$ and $D^{n,1}$.
@@ 3.2 The Hybrid Method
$\mathrm{Adv}(D_{\mathrm{unif}},D_{\mathrm{bias}})$ $=\mathrm{Adv}\left(D^{n,1},D^{1,n}\right)$
@@
In order to prove Theorem˜3.3, it suffices to consider its corresponding two-player communication problem, called Hidden Sign Problem (HSP).
Definition 3.4 (Hidden Sign Problem, I).
  • Case ($\perp$): Draw $Y^{a},Y^{b}\sim\{\pm 1\}^{m}$ independently.
  • Case (=): Draw $Y^{a}=Y^{b}\sim\{\pm 1\}^{m}$.
Once $Y^{a},Y^{b}$ are drawn, Alice (resp. Bob) receives a sequence of $n/2$ samples from $P_{Y^{a}}$ (resp. $P_{Y^{b}}$). Their goal is to distinguish between the two cases.
We will apply a Poissonization trick and analyze the following variant of the Hidden Sign Problem.
Definition 3.5 (Hidden Sign Problem, II).
@@
A reduction.
An equivalent way to state the input distributions for Version II of the Hidden Sign Problem is that each player will first draw an integer $n_{a}$ (resp. $n_{b}$) from $\mathrm{Poi}(n)$, and then draw $n_{a}$ ($n_{b}$) samples from $P_{Y^{a}}$ (resp. $P_{Y^{b}}$). It is a standard fact that $\Pr[\mathrm{Poi}(n)\geq\frac{n}{2}]\geq\frac{1}{2}$. As a consequence, with probability at least $\frac{1}{4}$, we have $\min(n_{a},n_{b})\geq n$. Thus, if we had a $C$-bit communication protocol that solves Version I of HSP with advantage $\xi$, we could then design a $(C+2)$-bit protocol that solves Version II with advantage at least $\frac{\xi}{4}$ (both players use $2$ bits to check that $\min(n_{a},n_{b})\geq n$, and run the assumed protocol with the first $n$ samples if the check passes).
Given this reduction, the rest of the paper will focus on proving a communication lower bound for the Hidden Sign Problem (Version II). That is, we will prove the following theorem.
@@
For every $\epsilon,m$, $n\leq O(m/\epsilon^{2})$, and $C\geq\log m$, it holds that that any $C$-bit protocol for the Hidden Sign Problem (II) achieves a distinguishing advantage of at most $O\left(\log^{2}(m)\cdot\frac{n}{m}\cdot C\cdot\epsilon^{2}\right)$.
We note that Theorem˜3.6 implies Theorem˜3.3, and hence the main result of our paper.
4 Proof of Theorem˜3.6
In this section, we prove Theorem˜3.6. This is achieved via a series of reductions.
@@ 4.1 The Hidden Index Problem
Finally, in order to analyze the game from Definition˜3.5, we will have to work with a variant of Noisy HIP where we observe not bits but some other forms of signal emitted from two similar sources.
@@ 4.3 Concluding the Proof
We are now ready to analyze the game from Definition˜3.5. We state the following theorem.